# RadOnc-GPT: An Autonomous LLM Agent for Real-Time Patient Outcomes Labeling at Scale

### v1

@@ Abstract
Purpose: Manual labeling limits the scale, accuracy, and timeliness of patient outcomes research in radiation oncology. We present RadOnc-GPT, an autonomous large language model (LLM)–based agent capable of independently retrieving patient-specific information, iteratively assessing evidence, and returning structured outcomes. Our evaluation explicitly validates RadOnc-GPT across two clearly defined tiers of increasing complexity: (1) a structured quality assurance (QA) tier, assessing the accurate retrieval of demographic and radiotherapy treatment plan details, followed by (2) a complex clinical outcomes labeling tier involving determination of mandibular osteoradionecrosis (ORN) in head-and-neck (H&N) cancer patients and detection of cancer recurrence in independent prostate and head-and-neck cancer cohorts requiring combined interpretation of structured and unstructured patient data. The QA tier establishes foundational trust in structured-data retrieval, a critical prerequisite for successful complex clinical outcome labeling.
@@
Methods: A wrapper program called LLM Task Streaming sequentially supplied patient IDs and specific task instructions to RadOnc-GPT. Initially, RadOnc-GPT completed the structured QA task using demographic and radiotherapy treatment plan course data exclusively. In the second tier, RadOnc-GPT autonomously combined structured electrical health record (EHR) data with unstructured clinical notes, radiology, and pathology reports to determine clinical outcomes. Ground-truth labels for ORN and cancer recurrence were retrospectively generated by expert radiation oncologists. Discrepancies between RadOnc-GPT’s outputs and ground-truth labels underwent independent adjudication by independent radiation oncologists, classifying each discrepancy into one of three categories: model error, ground-truth error, or indeterminate.
@@
Results: Across 895 unique patients, RadOnc-GPT matched all six demographic fields for all 500 patients in the QA tier (100%) and accurately reproduced radiation-course counts in 497/500 (99.4%), demonstrating robust retrieval fidelity. In the complex patient outcomes labeling tier, initial accuracy improved significantly post-adjudication: for ORN determination (233 patients) accuracy rose from 84.5% to 95.2%, prostate cancer recurrence detection (80 patients) improved from 92.5% to 95.0%, and H&N recurrence detection (82 patients) improved from 92.7% to 96.3%. Among 48 initial discrepancies, adjudication revealed 30 (63%) previously unrecognized ground-truth errors, 13 genuine model errors, and 5 indeterminate cases.
@@
Conclusions: RadOnc-GPT reliably retrieves foundational structured data and subsequently generalizes complex clinical outcome labeling tasks effectively, notably using a single cancer recurrence detection prompt across multiple disease sites. Its high recall performance minimizes clinically critical false negatives, while its dual capacity as labeler and auditor significantly enhances registry data integrity. Thus, RadOnc-GPT enables scalable, trustworthy, and real-time curation of radiation-oncology research datasets.
@@ 1 Introduction
Large-language models (LLMs) have progressed rapidly from general-purpose chatbots to highly specialized systems capable of matching or surpassing human experts across multiple clinical tasks[1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Specifically, in radiation oncology GPT-4 has attained expert-level performance on radiation-oncology physics exams [11, 12], broad clinical knowledge assessments [13, 14, 15], clinical task management such as in-basket response [16], automated drafting of payer-appeal documents [17], and harmonizing naming conventions of anatomical structures [18]. In addition, recently open source LLMs have been fine-tuned to improve the performance of important clinical tasks in radiation oncology [19]. Further improvements have been demonstrated through retrieval-augmented generation (RAG) approaches, which significantly enhance accuracy in clinical trial screening applications [20, 21, 22, 23, 24].
@@
Parallel studies in oncology-specific natural language processing (NLP) have demonstrated the successful extraction of detailed clinical information such as pathology findings [25], computed tomography (CT) imaging phenotypes [26], key entities from discharge letters [27], toxicity assessments [28], patient comorbidities [29], outcomes from lung-ablation procedures [30], and patient-reported outcomes [31]. Recent multi-modal models that integrate imaging and textual data have further expanded capabilities to tasks like tumor contouring [32] and survival prediction [33]. However, a significant limitation of most existing NLP approaches and registry curation efforts is their inherent assumption that institutional registry labels represent reliable gold standards. Recent audits, such as a multi-year review of an academic head-and-neck cancer registry, suggest otherwise, revealing significant latent errors, with up to half of true recurrence events potentially missing from automated flags [34]. Correcting these errors via traditional manual review would require substantial staffing resources, highlighting the urgent need for more efficient, scalable methods for auditing and curating clinical outcomes data.
@@
To address these critical limitations, we developed RadOnc-GPT, a GPT-4o–based autonomous agent capable of retrieving structured and unstructured clinical data directly from institutional databases, iteratively assessing the evidence, and synthesizing structured clinical outcomes. In order to evaluate RadOnc-GPT, we designed a two-tier strategy differing by their level of complexity: first, a relatively simple structured data retrieval task aimed at establishing the agent’s structured information retrieval capability. Because this task involves reproducing structured data, manual review is not required and therefore makes this evaluation step a good candidate for quality assurance purposes. After demonstrating reliable structured-data reproduction, we increase the complexity of the tasks, evaluating RadOnc-GPT’s ability to mix structured and unstructured patient data towards performing the task. These high-complexity tasks included determination of mandibular osteoradionecrosis (ORN) in head-and-neck cancer patients and cancer recurrence detection in separate prostate and head-and-neck cancer patient cohorts. A noteworthy design feature of our evaluation was using an identical cancer recurrence detection method (identical input prompt) across two disease sites—prostate cancer and head-and-neck cancer—testing RadOnc-GPT’s ability to generalize clinical reasoning across disease sites.
@@
Answering these questions clearly and rigorously establishes autonomous LLM agents as a potentially transformative solution for near-real-time, scalable and accurate clinical outcomes labeling.
@@ 2.1 RadOnc-GPT System Design
Figure 1 illustrates the architecture of RadOnc-GPT, which integrates both internal and external data resources. Internal connections include Mayo Clinic’s radiation oncology–specific database system (Minnesota, Arizona, Florida), Aria (Varian Medical Systems, Palo Alto, CA), and enterprise electronic health record (EHR) database systems interfacing with Epic (Epic Systems, Verona, WI). RadOnc-GPT further connects with public data resources including PubMed, ClinicalTrials.gov, and the National Cancer Institute (NCI) Common Terminology Criteria for Adverse Events (CTCAE), each connected via their public APIs. RadOnc-GPT is designed as an LLM-based agent, defined here as a system capable of conducting multi-turn conversations while autonomously determining which functions to call and when to stop responding. The RadOnc-GPT system makes use of the many data sources by calling upon whitelisted functions that retrieve, process, and return data to GPT-4o (OpenAI) or Gemini 2.5 Pro (Google). The specifications of these functions—including their inputs and outputs—are explicitly detailed within the system prompt so as to promote good decision making on the part of the LLM in deciding which functions to run and when to continue or stop responding.
@@ 2.2 Data Retrieval Functions
Table 1 summarizes the whitelisted functions available to RadOnc-GPT. When invoked, each function executes and returns outputs directly to the model. These outputs are designed to provide a coherent, structured representation of the underlying data, thereby maximizing the LLM’s ability to interpret and reason over the information. If no relevant data were found by a particular data retrieval function, the function would explicitly indicate that no data could be found for the given inputs and patient.
Importantly, RadOnc-GPT does not employ conventional retrieval-augmented generation (RAG). In RAG, embedded documents are matched to queries via vector similarity and delivered back to the model as text chunks—a strategy developed to surface relevant passages from vast, poorly organized unstructured corpora such as the internet, which often exceed model context limits. By contrast, RadOnc-GPT leverages the fact that patient data within Epic is unstructured yet systematically organized and indexed. Clinical notes, for example, are timestamped, provider-signed, and tagged with metadata such as medical specialty and treating department, making them straightforward to retrieve without generic RAG pipelines. As LLMs continue to support dramatically larger contexts (e.g., GPT-4o accommodates up to 1 million tokens) while inference costs per token decline (\$2.50 per 1 million input tokens for OpenAI’s standard GPT-4o now), the necessity of general-purpose RAG approaches, which introduce unnecessary complications to the system, further diminishes in settings like RadOnc-GPT, where targeted, well-structured data retrieval suffices. For instance, when managing a prostate cancer patient, RadOnc-GPT retrieves all radiation oncology and urology clinical notes relevant to that patient.
To avoid overwhelming the model with unnecessarily lengthy inputs, RadOnc-GPT employs a large set of highly specific retrieval functions. The guiding principle is that greater function granularity reduces the volume of data returned on a per-function basis. In rare cases where retrieved content exceeds the maximum allowable context length, a controlled pruning strategy is applied. The system prompt remains intact, while prior messages are selectively truncated. Because all retrieval functions return results in reverse chronological order (newest first, oldest last), pruning proceeds by removing fixed-size batches of words from the ends of the oldest messages. After each pass, the token count is recalculated; if the context remains too long, pruning continues until the message reaches a minimum threshold, at which point the process shifts to the next oldest message. Messages already at or below a skip threshold are preserved until all messages fall below that limit; if the conversation still exceeds the token budget, the threshold is lowered stepwise and pruning resumes. For this study, parameters were set as follows: maximum conversation history = 95,000 tokens, words removed per pass = 250, minimum threshold = 10,000 tokens, and threshold decrement = 2,500 tokens.
@@ 2.3 Two-Tier Evaluation Scaffold
RadOnc-GPT’s capability to extract patient outcomes was evaluated in two distinct tiers of escalating complexity (see Figure 2). This two-tiered approach establishes foundational accuracy in basic retrieval tasks first, followed by more clinically relevant tasks of clinical outcomes labeling:
@@
  • Tier 2 (Patient outcomes labeling from structured and unstructured patient data): In the second evaluation, RadOnc-GPT was assessed on three clinically complex outcome-labeling tasks: (1) detection of mandibular osteoradionecrosis (ORN) in patients with head-and-neck cancer following radiation therapy, (2) detection of cancer recurrence in prostate cancer patients after radiation therapy, and (3) detection of recurrence in head-and-neck cancer patients after surgery. These tasks required RadOnc-GPT to autonomously synthesize both structured data (e.g., demographics, treatment planning details, diagnosis codes) and unstructured data (e.g., clinical notes, radiology reports, pathology reports). For recurrence detection, the same prompt design was applied across prostate and head-and-neck cohorts, allowing evaluation of RadOnc-GPT’s ability to generalize across disease sites and treatment modalities. The outcomes labels generated by RadOnc-GPT were compared with ground truth labels that were acquired through careful manual review by experienced physicians. Our study considers that ground truth may not be accurate. Each discrepancy between RadOnc-GPT and the ground truth was therefore adjudicated by an independent physician for a final result.
Figure 2: Tiered difficulty of tasks. Two-Tier evaluation scaffold where tier 1 is relatively simple, structured data-only extraction task, and tier 2 is a difficult task requiring structured and unstructured data. Tier 1 evaluation does not require manual review; Tier 2 does.
@@ 2.4 LLM Task Orchestration
To enable large-scale evaluation, we developed an external orchestration framework called LLM Task Streaming, designed to allow for processing a population of patients with RadOnc-GPT. See Figure 3. This system sequentially supplied RadOnc-GPT with patient identifiers from each cohort along with task-specific instructions (prompts). The model’s outputs were returned as structured JSON objects, which were automatically parsed and consolidated into cohort-level CSV files for downstream analysis and physician adjudication. The framework supported parallel processing, handling up to four patients concurrently. In practice, overall throughput was constrained primarily by the inference speed of GPT-4o, with each patient’s data typically processed within 10–20 seconds depending on the volume of retrieved records.
@@ 2.5 Prompts
To execute a given task, RadOnc-GPT requires two inputs: the patient identifier and a task-specific prompt. The full set of prompts used for each task is provided in the Appendix. In general, all prompts followed a consistent design strategy: they specified the overall task, outlined which patient data should be retrieved, and included task-specific instructions. These task-specific components could incorporate definitions, clinical criteria, or reasoning strategies (e.g., chain-of-thought guidance) to ensure accurate and consistent outputs. The process of developing the prompts involved frequent observation and feedback from experienced physicians.
@@ 2.6 Cohorts and Ground Truth Labeling
ORN detection (head-and-neck) 233 34 199
Prostate recurrence detection 80 40 40
Head-and-neck recurrence detection 82 48 34
@@
Selection: Head-and-neck cancer patients were included if they met prespecified criteria: adults (>>19 years) treated between April 2013 and August 2019 with fraction sizes of 1.2–2.2 Gy and a total dose >>30 Gy, who subsequently developed mandibular osteoradionecrosis (ORN) after their initial radiotherapy course. Patients with recurrent cancer were excluded.
Ground truth: ORN status was established using CTCAE v4 osteonecrosis-of-the-jaw criteria and graded according to the Marx system. Positivity required documentation of ORN within the irradiated volume, supported by clinical notes, imaging, and lesion photographs. Alternative etiologies (e.g., infection, surgery, dental extraction, lymphoma, or disease outside the radiation field) were excluded.
Head-and-neck cancer recurrence cohort.
Selection: Head-and-neck cancer patients with prior radiation therapy.
Ground truth: Recurrence status was determined through chart review by an experienced clinical data abstractor (¿$15$ years), using predefined criteria established by a surgical oncologist. Ambiguities were resolved through weekly review sessions with the oncologist. All head-and-neck primary sites and histologies were eligible for inclusion.
@@
Selection: Prostate cancer patients treated with definitive, curative-intent radiotherapy.
Ground truth: Recurrence was defined as any documented post-treatment progression of prostate cancer and was labeled if either criterion was met: (1) biochemical recurrence (PSA nadir + 2 ng/mL) or a clinician-determined, intervention-triggering PSA rise; and/or (2) clinical recurrence—local, regional, distant nodal, or metastatic progression confirmed by imaging or biopsy. Only prostate cancer recurrences were considered when other primary cancers were present.

### v2

@@ Abstract
Purpose: Manual labeling limits the scale, accuracy, and timeliness of patient outcomes research in radiation oncology. We present RadOnc-GPT, an autonomous large language model (LLM)–based agent capable of independently retrieving patient-specific information, iteratively assessing evidence, and returning structured outcomes. Our evaluation explicitly validates RadOnc-GPT across two clearly defined tiers of increasing complexity: (1) a structured quality assurance (QA) tier, assessing the accurate retrieval of demographic and radiotherapy treatment plan details, followed by (2) a complex clinical outcomes labeling tier involving determination of mandibular osteoradionecrosis (ORN) in head-and-neck cancer (HNC) patients and detection of cancer recurrence in independent prostate and HNC cohorts requiring combined interpretation of structured and unstructured patient data. The QA tier establishes foundational trust in structured-data retrieval, a critical prerequisite for successful complex clinical outcome labeling.
@@
Methods: A wrapper program called LLM Task Streaming sequentially supplied patient IDs and specific task instructions to RadOnc-GPT. Initially, RadOnc-GPT completed the structured QA task using demographic and radiotherapy treatment plan course data exclusively. In the second tier, RadOnc-GPT autonomously combined structured electrical health record (EHR) data with unstructured clinical notes, radiology, and pathology reports to determine clinical outcomes. Ground-truth labels for ORN and cancer recurrence were retrospectively generated by expert radiation oncologists. Discrepancies between RadOnc-GPT’s outputs and ground-truth labels underwent adjudication by independent radiation oncologists, classifying each discrepancy into one of three categories: model error, ground-truth error, or indeterminate.
@@
Results: Across 895 unique patients, RadOnc-GPT matched all six demographic fields for all 500 patients in the QA tier (100%) and accurately reproduced radiation-course counts in 497/500 (99.4%), demonstrating robust retrieval fidelity. In the complex patient outcomes labeling tier, initial accuracy improved significantly post-adjudication: for ORN determination (233 patients) accuracy rose from 84.5% to 95.2%, prostate cancer recurrence detection (80 patients) improved from 92.5% to 95.0%, and HNC recurrence detection (82 patients) improved from 92.7% to 96.3%. Among 48 initial discrepancies, adjudication revealed 30 (63%) previously unrecognized ground-truth errors, 13 genuine model errors, and 5 indeterminate cases.
@@
Conclusions: RadOnc-GPT reliably retrieves foundational structured data and subsequently generalizes complex clinical outcome labeling tasks, notably using a single cancer recurrence detection prompt across multiple disease sites. Its high recall performance minimizes clinically critical false negatives, while its dual capacity as labeler and auditor significantly enhances registry data integrity. Thus, RadOnc-GPT enables scalable, trustworthy, and real-time curation of radiation-oncology research datasets.
@@ 1 Introduction
Large-language models (LLMs) have progressed rapidly from general-purpose chatbots to highly specialized systems capable of matching or surpassing human experts across multiple clinical tasks[RN1787, RN1790, RN1169, RN1789, RN1559, RN1553, RN1389, RN1788, RN1445, RN1793]. Specifically, in radiation oncology GPT-4 has attained expert-level performance on radiation-oncology physics exams [holmes2023evaluating, wang2024recent], broad clinical knowledge assessments [ramadan2025evaluating, longwell2024performance, dennstadt2024exploring], clinical task management such as in-basket response [hao2025retrospective], automated drafting of payer-appeal documents [kiser2024large], and harmonizing naming conventions of anatomical structures [holmes2024benchmarking]. In addition, recently open source LLMs have been fine-tuned to improve the performance of important clinical tasks in radiation oncology [RN1791]. Further improvements have been demonstrated through retrieval-augmented generation (RAG) approaches, which significantly enhance accuracy in clinical trial screening applications [doi:10.1056/AIoa2400181, PMID:37576126, RYBINSKI2024104734, ferber2024endtoendclinicaltrialmatching, gupta2024prismpatientrecordsinterpretation].
@@
Parallel studies in oncology-specific natural language processing (NLP) have demonstrated the successful extraction of detailed clinical information such as pathology findings [Geevarghesejcp-2024-209669], computed tomography (CT) imaging phenotypes [doi:10.1148/radiol.231362], key entities from discharge letters [SIEPMANN2025100378], toxicity assessments [khanmohammadi2024novel], patient comorbidities [info:doi/10.2196/58457], outcomes from lung-ablation procedures [GEEVARGHESE2024], and patient-reported outcomes [doi:10.1200/CCI.23.00258]. Recent multi-modal models that integrate imaging and textual data have further expanded capabilities to tasks like tumor contouring [oh2024llm] and survival prediction [cancers16132402]. However, a significant limitation of most existing NLP approaches and registry curation efforts is their inherent assumption that institutional registry labels represent reliable gold standards. Recent audits, such as a multi-year review of an academic HNC registry, suggest otherwise, revealing significant latent errors, with up to half of true recurrence events potentially missing from automated flags [SUTTON2024225]. Correcting these errors via traditional manual review requires substantial staffing resources, highlighting the need for more efficient, scalable methods for auditing and curating clinical outcomes data.
@@
To address these critical limitations, we developed RadOnc-GPT, a GPT-4o–based autonomous agent capable of retrieving structured and unstructured clinical data directly from institutional databases, iteratively assessing the evidence, and synthesizing structured clinical outcomes. In order to evaluate RadOnc-GPT, we designed a two-tier strategy differing by level of complexity: first, a relatively simple structured data retrieval task aimed at establishing the agent’s structured information retrieval capability. Because this task involves reproducing structured data, manual review is not required and therefore makes this evaluation step a good candidate for quality assurance purposes. After demonstrating reliable structured-data reproduction, we increase the complexity of the tasks, evaluating RadOnc-GPT’s ability to mix structured and unstructured patient data towards performing the task. These high-complexity tasks included determination of mandibular osteoradionecrosis (ORN) in HNC patients and cancer recurrence detection in separate prostate and HNC patient cohorts. A noteworthy design feature of our evaluation was using an identical cancer recurrence detection method (identical input prompt) across two disease sites, prostate cancer and HNC, testing RadOnc-GPT’s ability to generalize clinical reasoning across disease sites.
@@
Answering these questions clearly and rigorously establishes autonomous LLM agents as a potentially transformative solution for near-real-time clinical outcomes labeling at scale.
@@ 2.1 RadOnc-GPT System Design
Figure 1 illustrates the architecture of RadOnc-GPT, which integrates both internal and external data resources. Internal connections include Mayo Clinic’s radiation oncology–specific database system (Minnesota, Arizona, Florida), Aria (Varian Medical Systems, Palo Alto, CA), and enterprise electronic health record (EHR) database systems interfacing with Epic (Epic Systems, Verona, WI). RadOnc-GPT further connects with public data resources including PubMed, ClinicalTrials.gov, and the National Cancer Institute (NCI) Common Terminology Criteria for Adverse Events (CTCAE), each connected via their public APIs. RadOnc-GPT is designed as an LLM-based agent, defined here as a system capable of conducting multi-turn conversations while autonomously determining which functions to call and when to stop responding. The RadOnc-GPT system makes use of the many data sources by calling upon whitelisted functions that retrieve, process, and return data to GPT-4o (OpenAI). The specifications of these functions—including their inputs and outputs—are explicitly detailed within the system prompt so as to promote good decision making on the part of the LLM in deciding which functions to run and when to continue or stop responding.
@@ 2.2 Data Retrieval Functions
Table 1 summarizes the whitelisted functions available to RadOnc-GPT. When invoked, each function executes and returns outputs directly to the model. These outputs are designed to provide a coherent, structured representation of the underlying data, thereby maximizing the LLM’s ability to interpret and reason over the information. If no relevant data were found by a particular data retrieval function, each function explicitly indicates that no data could be found for the given inputs.
Notably, RadOnc-GPT does not employ conventional retrieval-augmented generation (RAG). In RAG, embedded documents are matched to queries via vector similarity and delivered back to the model as text chunks. This strategy was developed to surface relevant passages from vast, poorly organized, unstructured corpora such as the internet or a library, which often greatly exceed model context limits. By contrast, RadOnc-GPT leverages the fact that patient data within Epic is unstructured yet systematically organized and indexed. Clinical notes, for example, are timestamped, provider-signed, and tagged with metadata such as medical specialty, making them straightforward to retrieve without generic RAG pipelines. As LLMs continue to support increasingly larger contexts while inference costs per token decline (\$2.50 per 1 million input tokens for OpenAI’s standard GPT-4o as of this report), the necessity of general-purpose RAG approaches further diminishes in medical settings such as RadOnc-GPT operates, where targeted, well-structured data retrieval suffices.
To avoid overwhelming the model with unnecessarily lengthy inputs, RadOnc-GPT employs a large set of highly specific retrieval functions. The guiding principle is that greater function granularity reduces the volume of data returned on a per-function basis. In rare cases where retrieved content exceeds the maximum allowable context length, a controlled pruning strategy is applied. The system prompt remains intact, while prior messages are selectively truncated. Because all retrieval functions return results in reverse chronological order (newest first, oldest last), pruning proceeds by removing fixed-size batches of words from the ends of the oldest messages. After each pass, the token count is recalculated; if the context remains too long, pruning continues until the message reaches a minimum threshold, at which point the process shifts to the next oldest message in the conversation. If all the messages have been pruned to the threshold and the conversation still exceeds the token budget, the threshold is lowered stepwise and pruning resumes. For this study, parameters were set as follows: maximum conversation history = 95,000 tokens, words removed per pass = 250, minimum threshold = 10,000 tokens, and threshold decrement = 2,500 tokens.
@@ 2.3 Two-Tier Evaluation Scaffold
RadOnc-GPT’s capability to extract patient outcomes was evaluated in two distinct tiers of escalating complexity. This two-tiered approach establishes foundational accuracy in basic retrieval tasks first, followed by more clinically relevant tasks of clinical outcomes labeling:
@@
  • Tier 2 (Patient outcomes labeling from structured and unstructured patient data): In the second evaluation, RadOnc-GPT was assessed on three clinically complex outcome-labeling tasks: (1) detection of mandibular osteoradionecrosis (ORN) in patients with HNC following radiation therapy, (2) detection of cancer recurrence in prostate cancer patients after radiation therapy, and (3) detection of cancer recurrence in HNC patients after surgery. These tasks required RadOnc-GPT to autonomously synthesize both structured data (e.g., demographics, treatment planning details, diagnosis codes) and unstructured data (e.g., clinical notes, radiology reports, pathology reports). For cancer recurrence detection, the same prompt design was applied across prostate cancer and HNC cohorts, allowing evaluation of RadOnc-GPT’s ability to generalize across disease sites and treatment modalities. The outcomes labels generated by RadOnc-GPT were compared with ground truth labels that were acquired through careful manual review by experienced physicians. Our study considers that ground truth may not be accurate. Each discrepancy between RadOnc-GPT and the ground truth was therefore adjudicated by an independent physician for a final result.
@@ 2.4 LLM Task Orchestration
To enable large-scale evaluation, we developed an external orchestration framework called LLM Task Streaming, designed to allow for processing a population of patients with RadOnc-GPT. See Figure 2. This system sequentially supplied RadOnc-GPT with patient IDs from each cohort along with task-specific instructions (prompts). The model’s outputs were returned as structured JSON objects, which were automatically parsed and consolidated into cohort-level CSV files for downstream analysis and physician adjudication. The framework supports parallel processing of patients, only limited by the artificial rate-limit associated with the account. In practice, data was processed in about 10–30 seconds per-patient depending on the volume of retrieved records.
@@ 2.5 Prompts
To execute a given task, RadOnc-GPT requires two inputs: the patient ID and a task-specific prompt. The full set of prompts used for each task are provided in the Appendix. In general, all prompts followed a consistent design strategy, specifying the overall task, data retrieval instructions, and task-specific instructions. These task-specific components could incorporate definitions, clinical criteria, or reasoning strategies (e.g., chain-of-thought guidance) to ensure accurate and consistent outputs. The process of developing the prompts involved frequent observation and feedback from experienced physicians.
@@ 2.6 Cohorts and Ground Truth Labeling
ORN detection (HNC) 233 34 199
Recurrence detection (Prost.) 80 40 40
Recurrence detection (HNC) 82 48 34
@@
HNC patients were included according to the following criteria: adults (>19>19 years) treated between April 2013 and August 2019 with fraction sizes of 1.2–2.2 Gy and a total dose >>30 Gy, who subsequently developed mandibular osteoradionecrosis (ORN) after their initial radiotherapy course. Patients with recurrent cancer were excluded. ORN status was established using CTCAE v4 osteonecrosis-of-the-jaw criteria and graded according to the Marx system. Positivity required documentation of ORN within the irradiated volume, supported by clinical notes, imaging, and lesion photographs. Alternative etiologies (e.g., infection, surgery, dental extraction, lymphoma, or disease outside the radiation field) were excluded.
HNC recurrence cohort.
HNC patients (oropharynx, squamous cell carcinoma, prior radiation therapy) were selected from the Department of Otolaryngology Oropharynx RedCap registry. Cancer recurrence status was determined through chart review by an experienced clinical data abstractor (>15>15 years), supervised by a senior head-and-neck surgeon and a senior surgical oncologist. Ambiguities were resolved through weekly review sessions with the oncologist.
@@
Select prostate cancer patients were treated with definitive, curative-intent radiotherapy. Cancer recurrence was defined as any documented post-treatment progression of prostate cancer and was labeled if either criterion was met: (1) biochemical recurrence (PSA nadir + 2 ng/mL) or a clinician-determined, intervention-triggering PSA rise; and/or (2) clinical recurrence: local, regional, distant nodal, or metastatic progression confirmed by imaging or biopsy. Only prostate cancer recurrences were considered when other primary cancers were present.